OpenStudy (anonymous):

The design of a digital box camera maximizes the volume while keeping the sum of the dimensions at 6 inches. If the length must be 1.5 time the height, what should each dimension be? Hint: Let x represent one of the dimensions, and then define the other dimensions in terms of x.

OpenStudy (ddcamp):

Let x be the height, l be the length, and w be the width. Can you find an equation to relate x and l?

OpenStudy (anonymous):

idk I am extremely confused

OpenStudy (ddcamp):

We're told that the length is 1.5 times the height, so l=1.5x What else are we told about the different lengths that we can make an equation out of?

OpenStudy (anonymous):

the sum has to be 6 inches???

OpenStudy (ddcamp):

Yes. Using x, l, and w, what would that equation be?

OpenStudy (anonymous):

idk....

OpenStudy (ddcamp):

The sum of the dimensions is length+width+height. Since we know this has to add up to 6, we have: x+l+w=6

OpenStudy (anonymous):

ok

OpenStudy (ddcamp):

Since we know that l=1.5x, we can say that x+1.5x+w=6, then solve for w in terms of x.

OpenStudy (anonymous):

how would you do that?

OpenStudy (ddcamp):

\[x+1.5x+w = 6 \\ 2.5x+w=6 \\ w=6-2.5x\]

OpenStudy (ddcamp):

Now we have length, width, and height all in terms of one variable. What would the volume be? l= 1.5x w= 6-2.5x h= x

OpenStudy (anonymous):

6.5?

OpenStudy (ddcamp):

\[(1.5x)(6-2.5x)(x) \\ 1.5x^2(6-2.5x) \\ 9x^2 - 3.75x^3\] Do you know how to find the maximum of that?

OpenStudy (anonymous):

no sorry

OpenStudy (ddcamp):

To find a relative max/min of a function, take the derivative and find the zeroes.

OpenStudy (anonymous):

i have no idea what u r saying

OpenStudy (ddcamp):

What class is this for?

OpenStudy (anonymous):

math

OpenStudy (ddcamp):

What level math? (Algebra, Geometry, Calculus?)

OpenStudy (anonymous):

Algebra

OpenStudy (anonymous):

pre algebra

OpenStudy (ddcamp):

Well, we got to our equation for volume, and we're trying to find the largest possible volume. I would recommend graphing the equation, then see where the greatest volume is. Other than that, I don't know how to solve this using pre-algebra math.